Visit with a company of your interest (You have two to turn in the information )

Your Task:

1. Collect the following Information

a. Age of the employees

b. Salary of each employee

c. highest degree possible

d. Gender of each employee

Example: for one employee

Employee number Age Salary Degree Gender
1. 32 $35,000 Master Female

2. Stem-and-Leaf Display for age of the employee

3. Descriptive statistics for the salary ( Include the mean, standard deviation, sample size , minimum and maximum.)

4. Histogram between degree and Salary

5. What is probability that employee is female?

6. What is the probability that employee makes above average salary (Look at the mean in number 2)?
Answer #1

Answer

Employee number Age Salary ($) Degree Gender
1 32 35000 master female
2 31 32000 master female
3 25 36000 master female
4 26 25000 master female
5 34 26000 bachler female
6 38 37000 master male
7 40 35000 master male
8 43 35000 master male
9 27 26000 bachler male
10 28 26000 bachler female
11 29 26000 bachler male
12 30 30000 master male
13 35 30000 master male
14 36 28000 bachler female
15 34 26000 bachler female
16 35 30000 master male
17 34 32000 master male
18 45 24000 bachler female
19 40 30000 master male
20 42 27000 bachler male
21 29 28000 master female
22 28 25000 bachler male
23 32 30000 master male
24 31 40000 master female
25 30 38000 master male
26 28 26000 bachler male
27 27 25000 bachler male
28 40 35000 master female
29 26 26000 bachler female
30 27 25000 master female

2.

stem leaf
2 5,6,6,7,7,7,8,8,8,9,9
3 0,0,1,1,2,2,4,4,4,5,5,8
4 0,0,0,2,3,5

3.

mean 29800
var 21406896.55
SD 4626.758752
sample size 30
min 25
max 45

4.

.

5.probability=14/30=0.46

6.probability=15/30=0.5